31) Assume the DJIA closed at 15,150 last night. The divisor is 0.123017848. Assume that 29 of the stocks in the index were unchanged today. One stock increased in value from $44.80 a share yesterday to $47.90 a share today. What is the DJIA index value at the close of trading today?

A) 15,175.20

B) 15,208.30

C) 15,365.60

D) 15,412.20

E) 15,524.10


32) An index has a market value of $695,200 at the beginning of the period and $741,900 at the end of the period. If you want the beginning index value to be 100, what is the ending index value?

A) 104.73

B) 105.42

C) 105.67

D) 105.89

E) 106.72


33) A tax free municipal bond provides a yield of 2.34%. What is the equivalent taxable yield on the bond given a 28% tax bracket?

A) 2.34%

B) 2.68%

C) 3.25%

D) 4.92%

E) 5.25%


34) Barnegat Light sold 200,000 shares in an initial public offering. The underwriter's explicit fees were $90,000. The offering price for the shares was $35, but immediately upon issue, the share price jumped to $43. What is the best estimate of the total cost to Barnegat Light of the equity issue?

A) $90,000

B) $1,290,000

C) $2,390,000

D) $1,690,000

E) $1,750,000
Answer #1

31. DJIA value last night =DJIA index *Divisor =15150*0.123017848=1863.7204
The new value =DJIA -Old Price+New Price =1863.7204-44.80+47.90 =1866.8204
New DJIA index value =New Value/Divisor =1866.8204/0.123017848=15175.20
Option a is correct option

32. Ending Value of Index =End Period Market Value/Beginning Period Market Value*100 =741900/695200*100 =106.72
(Option e is correct option)

33. equivalent taxable yield =Yield/(1-Tax rate) =2.34%/(1-28%) =3.25%(Oprion c is correct Option)